OpenStudy (anonymous):

What fraction is closest to 1/2 3/9, 3/8, 3/4

TheSmartOne (thesmartone):

What do you mean by closest?

OpenStudy (anonymous):

i would think it would be 3/8

OpenStudy (anonymous):

undead we might have a pedophile in chat

TheSmartOne (thesmartone):

1/2=.5 3/9=.33333333333333333 3/8=.375 3/4=.75

TheSmartOne (thesmartone):

@Micahc

OpenStudy (anonymous):

Ok thanks